# Correspondence between two gravitational lens equations in a static and spherically symmetric spacetime

### v2

@@ I Introduction
This paper is organized as follows. Section II discusses the unphysical branch in VE equation. Section III clarifies a relation between the improved VE equation and OB one. Section IV argues possible numerical differences when the deflection angle transformation is ignored. Throughout this paper, we focus on a static and spherically symmetric spacetime.
@@ Acknowledgements.
We thank Shwetketu Virbhadra for comments on the earlier version of our manuscript. We wish to thank an anonymous referee for helpful comments. We are grateful to Keita Takizawa and Marcus Werner for fruitful discussions. We thank Yuuiti Sendouda and Ryuichi Takahashi for useful conversations. This work was supported in part by Japan Science and Technology Agency (JST) SPRING, Grant Number, JPMJSP2152 (R.K.) and in part by Japan Society for the Promotion of Science (JSPS) Grant-in-Aid for Scientific Research, No. 20K03963 and 24K07009(H.A.).

### v3

@@ I Introduction
This paper is organized as follows. Section II discusses the unphysical branch in VE equation. Section III clarifies a relation between the improved VE equation and OB one. Section IV argues possible numerical differences when the deflection angle transformation is ignored. Appendix A shows that there exists no solution corresponding to a lensed image in the unphysical branch of VE equation. Throughout this paper, we focus on a static and spherically symmetric spacetime.
@@ Acknowledgements.
We thank Shwetketu Virbhadra for helpful comments on the earlier version of our manuscript. We wish to thank an anonymous referee for helpful comments. We are grateful to Keita Takizawa and Marcus Werner for fruitful discussions. We thank Yuuiti Sendouda and Ryuichi Takahashi for useful conversations. This work was supported in part by Japan Science and Technology Agency (JST) SPRING, Grant Number, JPMJSP2152 (R.K.) and in part by Japan Society for the Promotion of Science (JSPS) Grant-in-Aid for Scientific Research, No. 20K03963 and 24K07009(H.A.).
Appendix A On the minus branch of VE equation
This appendix shows that there exists no solution corresponding to a lensed image in the minus branch of VE equation
$\tan\theta=\frac{D_{S}(\tan\beta\tan\alpha_{VE}-1)-\sqrt{\cal D}}{2D_{L}\tan\alpha_{VE}},$ (21)
in a relativistic compact lens object. Here, a relativistic compact lens object has a photon sphere around the lens, as represented by the Schwarzschild lens. Therefore, the closest distance $r_{0}$ of a photon is small, namely $r_{0}\ll D_{S}$.
VE equation considers the observer and the source in asymptotic flat regions but not in the neighborhood of the lens. Hence, $(D_{L}D_{LS})/(D_{S})^{2}\ll 1$ is not allowed. VE equation assumes also a nearly coaligned configuration, namely a small $\tan\beta$. Hence, it suffices to assume $|\tan\beta|<1$ in this appendix.
For the convenience in the following calculations, we introduce $r\equiv D_{L}/D_{S}$, $s\equiv\tan\beta$ and $t\equiv(\tan\alpha_{VE})^{-1}$ to reexpress Eq. (21) as
$\tan\theta=\frac{s-f(t)}{2r},$ (22)
where $D_{LS}=(1-r)D_{S}$, and
$f(t)\equiv t+\frac{t}{|t|}\sqrt{\tilde{\cal D}},$ (23)
$\tilde{\cal D}\equiv t^{2}{\cal D}/(D_{S})^{2}=p+qt+t^{2}$ for $p\equiv s^{2}+4r(1-r)$ and $q\equiv-2s(1-2r)$. Note that $10<r<$ 1, $40<r(1-r)\leq 1/$ 4, and $4p>q^{2}$. Without loss of generality, we can choose $s\geq 0$. We shall examine separately (1) $t>0$ and (2) $t<0$ below.
(1) $t>0$ case. Then, $f(t)=t+\sqrt{p+qt+t^{2}}$, for which $df(t)/dt=(\sqrt{p+qt+t^{2}}+t+q/2)/\sqrt{p+qt+t^{2}}$. Here, $\sqrt{p+qt+t^{2}}+t+q/2>0$, because $4p>q^{2}$. Hence, $df(t)/dt>0$ for $t>0$. Hence, $f(t)>f(0)=\sqrt{p}$.
We thus obtain
$|\tan\theta|$ $=\frac{1}{2r}\left|s-f(t)\right|$
$=\frac{1}{2r}\left(f(t)-s\right)$
$>\frac{1}{2r}(\sqrt{p}-s),$ (24)
where $f(t)>\sqrt{p}>s$ is used in the second and third lines.
For $s<1$, $\sqrt{p}-s>\sqrt{1+4r(1-r)}-1=(4r(1-r))/(1+\sqrt{1+4r(1-r)})\geq(4r(1-r))/(1+\sqrt{2})$, where $1+4r(1-r)\leq 2$ is used for the last inequality. By using this inequality for the last line in Eq. (24), $|\tan\theta|>(2/(1+\sqrt{2}))(D_{LS}/D_{S})$, which leads to the closest distance of a photon $r_{0}=|\tan\theta|\times D_{L}>(2/(1+\sqrt{2}))(D_{L}D_{LS}/D_{S})$. However, this inequality contradicts with $r_{0}\ll D_{S}$, because $(D_{L}D_{LS})/(D_{S})^{2}\ll 1$ is not allowed. For $t>0$, therefore, the minus branch has no solution corresponding to a lensed image.
(2) $t<0$ case. Then, $f(t)=t-\sqrt{p+qt+t^{2}}$, for which $df(t)/dt=(\sqrt{p+qt+t^{2}}-t-q/2)/\sqrt{p+qt+t^{2}}$. Here, $\sqrt{p+qt+t^{2}}-t-q/2>0$, because $4p>q^{2}$. Hence, $df(t)/dt>0$ for $t<0$. Hence, $f(t)<f(0)=-\sqrt{p}$.
We thus find
$|\tan\theta|$ $=\frac{1}{2r}\left|s-f(t)\right|$
$=\frac{1}{2r}(|f(t)|+s)$
$>\frac{1}{2r}(\sqrt{p}+s),$ (25)
where $f(t)<-\sqrt{p}<0$ is used in the second and third lines.
By using $\sqrt{p}+s\geq 2\sqrt{r(1-r)}$ for Eq. (25), we find $|\tan\theta|>\sqrt{D_{LS}/D_{L}}$, which leads to the closest distance of a photon $r_{0}=|\tan\theta|\times D_{L}>\sqrt{D_{L}D_{LS}}$. However, this inequality contradicts with $r_{0}\ll D_{S}$, because $(D_{L}D_{LS})/(D_{S})^{2}\ll 1$ is not allowed. For $t<0$ also, therefore, the minus branch has no solution corresponding to a lensed image.
In either case, the minus branch has no solution corresponding to a lensed image. Therefore, Eq. (21) is unphysical.